# Human-Object Interaction with Vision-Language Model Guided Relative Movement Dynamics

### v1

@@ Abstract
Human-Object Interaction (HOI) is vital for advancing simulation, animation, and robotics, enabling the generation of long-term, physically plausible motions in 3D environments. However, existing methods often fall short of achieving physics realism and supporting diverse types of interactions. To address these challenges, this paper introduces a unified Human-Object Interaction framework that provides unified control over interactions with static scenes and dynamic objects using language commands. The interactions between human and object parts can always be described as the continuous stable Relative Movement Dynamics (RMD) between human and object parts. By leveraging the world knowledge and scene perception capabilities of Vision-Language Models (VLMs), we translate language commands into RMD diagrams, which are used to guide goal-conditioned reinforcement learning for sequential interaction with objects. Our framework supports long-horizon interactions among dynamic, articulated, and static objects. To support the training and evaluation of our framework, we present a new dataset named Interplay, which includes multi-round task plans generated by VLMs, covering both static and dynamic HOI tasks. Extensive experiments demonstrate that our proposed framework can effectively handle a wide range of HOI tasks, showcasing its ability to maintain long-term, multi-round transitions. For more details, please refer to our project webpage: https://rmd-hoi.github.io/.
Figure 1: Our Relative Movement Dynamics (RMD) architecture enables dynamic object interaction and long-term multi-task completion based on VLM guidance. RMD enables the automatic construction of a unified reward function applicable to various reinforcement learning interaction tasks.
@@ 1 Introduction
Methods Dynamic Object Articulated Object Automated Reward Design Long-horizon Transition HOI Goal Definition Multitask
Kinematic NSM [30] -
SAMP [8] -
SceneDiffuser [11] -
OMOMO [14] -
Physics-Based InterPhys [9] One-to-One
PhysHOI [40] Multi-to-Multi
UniHSI [45] One-to-One
RMD (Ours) Multi-to-Multi
Table 1: Comparative Analysis of Key Features between RMD and Other Methods. The proposed framework enables a unified and automated design of goal states and reward functions. Additionally, the framework facilitates fine-grained modeling of spatio-temporal human-object interactions, ensuring natural and continuous interactions with static, dynamic, and articulated objects.
Human-Object Interaction (HOI) is pivotal for progress in fields like simulation, animation and robotics. The key research goal is to generate long-term, physically plausible human motions with corresponding object movements. Existing kinematic-based data-driven approaches [64, 24, 63, 62] often struggle with artifacts and lack of realism, as they depend on static or pre-defined object trajectories that fail to capture the nuanced, interactive dynamics between humans and moving objects.
Alternatively, some recent works employ physics engines [37, 21] to create interactive environments where agents can learn interaction skills through goal-conditioned reinforcement learning [27, 56, 40, 41, 9]. These methods take advantage of the physical realism provided by simulations to train agents in a variety of tasks, resulting in more plausible motion sequences. However, existing works primarily focus on specific task designs (e.g., sitting [3, 57] or carrying [50, 33, 4, 7]), which introduces two main issues: 1) Various interaction tasks require carefully crafted, task-specific rewards designed by human experts. Due to the complexity of dynamic human-object interactions, these reward functions are extremely challenging to design [32]; 2) These task-specific training approaches require separate policy networks and struggle with long-horizon, multi-round transitions, as the sparse goal settings often leave agents in unrecoverable states for subsequent skill transitions.
Recent efforts, such as Eureka [20], have attempted to address the first issue by leveraging large language models to automate reward design and reveal latent reward structures beyond human intuition. However, the rewards generated in this manner frequently lack interpretability and do not align well with human perceptual cues, often resulting in unnatural, non-human-like motion. UniHSI [45] proposed a “chain of contacts” framework leveraging the power of Large Language Models. However, this framework focuses on “contact” as an instantaneous spatial relationship, neglecting the temporal dynamics with objects, thus making it unsuitable for handling dynamic and complex interaction tasks. Moreover, this point-goal-reaching formulation fails to fully account for the coordination of whole-body movements, resulting in suboptimal behaviors that lack the natural fluidity and human-like motion essential for realistic interactions.
Inspired by the concept of relative motion in classical mechanics [53], we propose a novel approach that models complex HOI as stable, relative movements between human and object parts. For example, when picking up a box, different body parts maintain consistent relative distances to specific points on the object, capturing not only moment-to-moment goals (e.g., approaching the box) but also transitional behaviors (e.g., moving the box with both hands). This relational constraint encapsulates the spatial, temporal, and combinatorial elements of dynamic HOI, allowing for smoother and more realistic interactions. We introduce a bipartite diagram named the Relative Movement Dynamics (RMD) to describe the spatial-temporal relationships between human parts and object parts during a sub-sequence. Benefiting from the rapid development of Vision-Language Models (VLMs) and their powerful capabilities in world knowledge, we utilize a VLM as a high-level Planner to perceive the surrounding scene and decouple text instructions into sub-sequences, automatically formatting sub-motions into the RMD bipartite diagram. Accordingly, RMD is then constructed as the target in the goal-conditioned reinforcement learning. In summary, our contributions include:
  • $\bullet$
@@
    We propose a general physics-based HOI synthesis framework leveraging the powerful world knowledge of VLMs, supporting long-term interactions with a diverse range of objects, including static, dynamic and articulated ones.
  • $\bullet$
    We introduce Relative Movement Dynamics (RMD), a spatial-temporal bipartite diagram designed to model fine-grained, complex dynamic interactions. RMD enables the automatic construction of a unified reward function applicable to various reinforcement learning interaction tasks.
  • $\bullet$
@@
    We create Interplay, a novel dataset including thousands of interaction plans with both static and dynamic interaction tasks. Extensive simulation results demonstrate the effectiveness of our approach.
@@ 2 Related Work
Figure 2: An overview of the concept of Relative Movement Dynamics and VLM-guided RMD Planner. RMD describes the spatial-temporal relationships between human parts and objects during a sub-sequence. We utilize a bipartite graph to formulate RMD, where each node corresponding to a human body part is uniquely connected to a node representing an object part. The edges in this graph encode the relative movement dynamics between the parts.
2.1 Human Motion Synthesis with Static Scenes
@@
In the field of character animation and motion synthesis, most research has focused on human motion interacting with static 3D scenes [16, 56, 51, 59]. Typically, researchers break down complex instructions into several key static poses within the scene [8, 39], and then use motion inpainting and optimization techniques to generate transitions between these key poses [38]. However, this approach often leads to mediocre and inconsistent motions due to the limited expressiveness of the intermediate sequences. Recently, diffusion-based methods like [10, 11, 34, 52, 61, 36] have achieved better results in human-scene synthesis. However, these data-driven approaches are limited to generating short-term motions due to constraints in the dataset, and the physical plausibility of the generated motions is not guaranteed. Apart from data-driven kinematic approaches, some studies have explored the problem within a reinforcement learning framework. For instance,  [60, 58] achieved long-term goal-oriented behaviors by leveraging motion primitives with specific reward designs. The UniHSI [45] framework decomposes instructions into a sequence of point-reaching control tasks under a unified formulation. However, it addresses only the spatial constraints of static scenes while overlooking the temporal dynamics of objects, rendering it inadequate for managing complex dynamic interactions.
@@
2.2 Kinematic-based Human Motion Synthesis with Dynamic Objects
@@
Research in human motion synthesis has increasingly focused on modeling interactions with dynamic objects [31, 12, 57, 54]. Diffusion-based frameworks [15, 14, 28, 25], guide motion generation with object trajectories but often lack realism due to predefined object paths that fail to account for physical plausibility. In contrast, InterDreamer [48] first generates human motion and subsequently uses a pre-trained world model to produce object trajectories, though this approach is limited by the simplicity of the world model, leading to inaccuracies in trajectory prediction. Some studies attempt to address these limitations by jointly modeling human-object interactions with supplementary guidance techniques, including relation intervention [42], contact prediction [5], and affordance estimation [25]. Truman [13], trained on a high-quality human-scene interaction (HSI) dataset, achieves dynamic stability through an auto-regressive diffusion model guided by action labels and waypoints. Nevertheless, kinematic models continue to face challenges such as penetration, sliding issues, and difficulties in generating long-term motion, requiring extensive annotations.
@@
2.3 Physics-based Human Motion Synthesis with Dynamic Objects
@@
To generate physically plausible motions, reinforcement learning methods have proven effective in training human-object interaction skills based on motion capture data [3, 22, 26, 47, 49, 43]. AMP [27] introduced an adversarial motion prior framework for realistic motion synthesis. InterPhys [9] further extended this framework by incorporating an HOI motion prior, achieving success in tasks such as sitting, lying, and carrying. Further advancements have led to successful applications in sports activities, including basketball [40, 41], tennis [55], soccer [46], skating [17], and football [19, 18]. Since these approaches often depend on heuristic designs and struggle to generalize to longer, multi-round scenarios. Our method seeks to address these limitations by introducing a Relative Movement Dynamics representation, enabling the model to capture both spatial and temporal dynamics effectively. This results in physically plausible and temporally coherent interactions, eliminating the need for manual annotations and enhancing realism in dynamic, long-term HOI tasks. The detailed comparisons of key features are listed in Tab. 1.
@@ 3 Method
In the following subsections, we first illustrate the concept of the proposed Relative Movement Dynamics (Sec. 3.1). Then, we demonstrate how we use VLM to perceive the scene context and decouple the text instruction into a plan in the form of RMD (Sec. 3.2). Finally, we elaborate on the construction of the RMD controller, which receives RMD plans as input and executes them step by step (Sec. 3.3).
3.1 Relative Movement Dynamics
Complex human-object interaction (HOI) tasks can be decomposed into several sub-sequences, where each human part maintains a consistent relative movement trend with its corresponding object part until the transition to the next sub-sequence. Therefore, we universally define the objective of the interaction plan as $D$, with the formulation as,
$\mathcal{D}=\{\mathcal{G}_{0},\mathcal{G}_{1},\ldots,\mathcal{G}_{m}\},$ (1)
where $\mathcal{G}_{i}$ denotes the $i$-th plan interaction step.$\mathcal{G}$ is composed with three parts: the target location for human root, the target location for object root and the planned Relative Movement Dynamics between human parts and object parts, with the formulation as,
$\mathcal{G}=\{\mathcal{T}_{H},\mathcal{B},\mathcal{T}_{O}\},$ (2)
$\mathcal{B}$ is a bipartite graph, where each node corresponding to a human body part is uniquely connected to a node representing an object part. The edges in this graph encode the relative movement dynamics between the parts. In this context, we assume that the human body consists of 15 parts, while the object consists of $k$ parts. In this way, we can formulate each $\mathcal{B}$ as,
=$\mathcal{B}=\{e_{ij}\mid p_{h_{i}}\in P_{H},p_{o_{j}}\in P_{O},w_{ij}\in\{0,1,2,3\}\},$ (3)
where $P_{H}$ and $P_{O}$ are the sets of human body parts and object parts, respectively. The set of human body parts is given by,
$P_{H}=\{p_{h_{1}},\dots,p_{h_{15}}\},$ (4)
The set of object parts is,
$P_{O}=\{p_{o_{1}},\dots,p_{o_{k}}\},$ (5)
Each edge $e_{ij}\in\mathcal{B}$ connects a human part $p_{h_{i}}\in P_{H}$ to an object part $p_{o_{j}}\in P_{O}$, with an edge weight $w_{ij}$ representing the state of relative movement dynamics between the two parts. The interpretation of the weights is as follows,
  • $w_{ij}=0$ indicates that $p_{h_{i}}$ and $p_{o_{j}}$ remain relatively stationary with respect to each other during this period.
  • $w_{ij}=1$ indicates that $p_{h_{i}}$ and $p_{o_{j}}$ are moving towards each other during this period.
  • $w_{ij}=2$ indicates that $p_{h_{i}}$ and $p_{o_{j}}$ are moving away from each other during this period.
  • $w_{ij}=3$ indicates that $p_{h_{i}}$ and $p_{o_{j}}$ do not maintain a stable relative motion trend during this period.
@@ 3.2 VLM-guided RMD Planner
We utilize GPT-4v [1] as our planner to infer the spatial-temporal relationships between humans and objects during interactions in specific contexts. The planner translates high-level language commands $\mathcal{L}$ into detailed plans $D$. As illustrated in Fig. 2, the RMD planner takes three main inputs: a language command $\mathcal{L}$, a top-view image representing the environmental context $\mathcal{C}$, and pre-defined prompts. Leveraging GPT-4v’s [1] vision capabilities, the RMD planner perceives the surrounding environment, identifies relevant objects, and infers their parts. To facilitate this process, we employ prompt engineering to ensure that GPT-4v [1] generates structured JSON documents, which can be easily encoded. By modifying both the instructions and the environment layout, we can generate diverse sequential interactions involving both static and dynamic objects. This approach allows us to create interaction datasets for evaluation, as detailed in the experiment section.
3.3 Policy Learning via RMD
Receiving the detailed plan output by the RMD Planner and background scenarios in the form of meshes and point clouds as input, the policy outputs specific torque for each joint to activate the agent and sequentially execute the plan. We begin by introducing the preliminaries of our method. Then, we describe the detailed formulation of the policy, including how the plan is translated into the goal state and how we construct the corresponding reward to encourage the agent to perform actions that align with the goal expectations.
Preliminary. Our physics-based animation framework is built upon goal-conditioned reinforcement learning. At each time step $t$, the agent samples an action from its policy $\pi(a_{t}\mid\mathbf{s}_{t},\mathbf{g}_{t})$ based on the current state $\mathbf{s}_{t}$ and the goal state $\mathbf{g}_{t}$. After executing the action, the environment transitions to the next state $\mathbf{s}_{t+1}$, and the agent receives a task reward $r^{G}(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1})$. To train control policies that enable characters to perform high-level tasks in a natural and life-like manner, we adopt the AMP framework [27], which introduces a discriminator to encourage the character to produce behaviors similar to those in the dataset by providing a style reward $r^{S}(\mathbf{s}_{t},\mathbf{s}_{t+1})$. Combining these two reward terms, the total reward at time step $t$ is,
$r_{t}=w^{G}r^{G}\left(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1}\right)+w^{S}r^{S}\left(\mathbf{s}_{t},\mathbf{s}_{t+1}\right).$ (6)
Note that AMP, as well as most previous methods, spends time designing the task-specific goal state and task-specific reward term. In contrast, our framework supports fully automated goal state design and switching, as well as a unified reward signal.
Goal state via RMD. At time step $t$, given the $i$-th plan interaction step $\mathcal{G}_{i}=\{\mathcal{T}_{H},\mathcal{B},\mathcal{T}_{O}\}$, we first encode each edge $e_{ij}$ from $\mathcal{B}$ by collecting the corresponding absolute position $\mathbf{p}_{h_{i}}^{p}$ and absolute velocity $\mathbf{p}_{h_{i}}^{v}$ of the human part joint $\mathbf{p}_{h_{i}}$, as well as the absolute position $\mathbf{p}_{o_{j}}^{p}$ and absolute velocity $\mathbf{p}_{o_{j}}^{v}$ of the nearest point in the object part point cloud $\mathbf{p}_{o_{j}}\in\mathbb{R}^{3}$ from the simulation environment. Next, we calculate the relative position $\tilde{\mathbf{p}}_{ij}$ and the relative velocity $\tilde{\mathbf{v}}_{ij}$ for each pair of points and transform these quantities into the agent’s coordinate system. For each edge, a one-hot encoding $w^{\prime}_{ij}$ is applied to represent $w_{ij}$. The one-hot vector $w^{\prime}_{ij}$ is then concatenated with the regularized relative position $\tilde{\mathbf{p}}_{ij}$ and relative velocity $\tilde{\mathbf{v}}_{ij}$ at time step $t$, yielding the rmd state representation $\mathbf{s}_{t}^{\text{rmd}}=\{\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},w^{\prime}_{ij}\}.$ Both $\mathcal{T}_{H}$ and $\mathcal{T}_{O}$ are in the form object(spatial relationship type). In our implementation, we first compute the object’s bounding box and get $s_{x}$, $s_{y}$, $s_{z}$, which correspond to the size of the object along the $x$-, $y$-, or $z$-axis, respectively. The position is then determined by shifting from the object center to the target according to the specific spatial relationship type. For instance, armchair(front) represents a position that is $(0.7s_{x},0,0)$ in the armchair’s local coordinate system. Next, we transform the positions of $\mathcal{T}_{H}$ and $\mathcal{T}_{O}$ into the agent’s coordinate system. This transformation guides both the agent’s and the object’s trajectories, yielding the destination state $\mathbf{d}_{t}=\{\mathbf{d}_{t}^{h},\mathbf{d}_{t}^{o}\}.$ To enhance the agent’s ability to perceive its surrounding local environment and avoid obstacles, a resolution-fixed height map $\mathbf{h}_{t}$ is incorporated, which records the heights of nearby object geometries aligned with the root’s facing direction [35, 45]. Furthermore, to enhance object perception, key parameters are extracted: the eight vertices of the bounding box $\mathbf{b}_{t}^{\text{ver}}$, the rotation angle $\mathbf{\theta}_{t}$, the linear velocity $\mathbf{v}_{t}$, and the angular velocity $\mathbf{w}_{t}$. These parameters define the object’s state $\mathbf{o}_{t}=\{\mathbf{b}_{t}^{\text{ver}},\mathbf{\theta}_{t},\mathbf{v}_{t},\mathbf{w}_{t}\}.$ Finally, combine $\mathbf{s}_{t}^{\text{rmd}}$, $\mathbf{d}_{t}$, $\mathbf{h}_{t}$, and $\mathbf{o}_{t}$, we obtain the complete set of goal state.

### v2

# Human-Object Interaction via Automatically Designed VLM-Guided Motion Policy

@@ Abstract
Human-object interaction (HOI) synthesis is crucial for applications in animation, simulation, and robotics. However, existing approaches either rely on expensive motion capture data or require manual reward engineering, limiting their scalability and generalizability. In this work, we introduce the first unified physics-based HOI framework that leverages Vision-Language Models (VLMs) to enable long-horizon interactions with diverse object types—including static, dynamic, and articulated objects. We introduce VLM-Guided Relative Movement Dynamics (RMD), a fine-grained spatio-temporal bipartite representation that automatically constructs goal states and reward functions for reinforcement learning. By encoding structured relationships between human and object parts, RMD enables VLMs to generate semantically grounded, interaction-aware motion guidance without manual reward tuning. To support our methodology, we present Interplay, a novel dataset with thousands of long-horizon static and dynamic interaction plans. Extensive experiments demonstrate that our framework outperforms existing methods in synthesizing natural, human-like motions across both simple single-task and complex multi-task scenarios. For more details, please refer to our project webpage: https://vlm-rmd.github.io/.
@@ 1 Introduction
Human-Object Interaction (HOI) understanding is fundamental to advancing embodied AI systems in simulation, animation, and robotics. Existing approaches to HOI learning generally fall into two categories. The first paradigm employs motion tracking policies to imitate reference demonstrations [1, 2, 3]. While these methods can reproduce specific interaction patterns, they face critical limitations: Their heavy reliance on high-quality motion capture data creates scalability bottlenecks, and their strict adherence to reference trajectories inherently restricts their ability to generate novel interactions beyond the training distribution.
The second paradigm adopts a task-centric perspective, developing specialized policies for individual interactions like sitting [4, 5] or carrying [6, 7, 8, 9]. However, these methods encounter two fundamental challenges. First, they require labor-intensive reward engineering by domain experts - a particularly demanding requirement given the complex dynamics and contact-rich nature of realistic HOI scenarios [10]. This manual reward design process inherently limits generalizability across interaction types, creating a need for automated objective formulation. Second, constrained by artificially designed, single-objective reward mechanisms, trained policies often overfit to specific behavioral patterns. Producing task-compliant but biomechanically unrealistic motions that violate natural human kinematics.
Recent efforts such as Eureka [11] and Grove [12] leverage large language models (LLMs) to automatically generate reward function code. However, they rely on iterative search paradigms that are often sample-inefficient and computationally costly. UniHSI [13] advances this line of work by unifying reward design through a “chain-of-contacts” abstraction, which models an interaction as a discrete sequence of contact events. While conceptually elegant, this formulation overlooks the motion dynamics of interactions and lacks the capacity to model whole-body coordination or accommodate dynamic objects. Consequently, it struggles to handle dynamic interactions and often produces jittery, suboptimal behaviors even in static interaction scenarios.
Figure 1: Our framework automatically constructs goal states and reward functions for a variety of interaction tasks in reinforcement learning. Guided by VLMs, the resulting motion policy enables physics-based characters to perform long-horizon interactions with both static and dynamic objects.
To address these limitations, we propose a physics-based HOI framework that leverages VLMs to automatically construct goal states and reward functions, guiding the motion policy to perform long-term interactions with diverse objects. Specifically, we aim to empower VLM to provide motion-level, interaction-aware guidance by leveraging its abilities in motion imagination and semantic reasoning. Drawing inspiration from the classical notion of relative motion [14], we introduce Relative Movement Dynamics (RMD), a structured representation that encodes the fine-grained spatio-temporal relationships between human parts and object parts across an interaction sub-sequence. For example, in the task of lifting a box, the relative spatial configuration between the hands and the box remains stable, providing a natural constraint that captures both discrete interaction goals (e.g., contact) and continuous dynamics (e.g., coordinated movement). RMD is designed to prompt VLMs to imagine interaction dynamics while effectively grounding their high-level reasoning in motion-level patterns. This formulation enables VLMs to move beyond symbolic planning and participate in dynamic skill composition, thereby supporting more expressive and generalizable HOI motion policy learning. Building on RMD, our framework automatically constructs unified goal states that support both static and dynamic interactions, along with automatically designed reward functions that guide motion patterns consistent with the plan.
Our framework supports long-horizon interactions with diverse object types. However, existing datasets typically focus either on static interactions [13] or on object rearrangement [6], and thus fail to cover this setting. To address this gap, we construct a new dataset, Interplay, which includes long-horizon static and dynamic interaction tasks across varied scene contexts. This dataset enables systematic evaluation of our framework in a long-horizon, multi-task scenario. Experimental results demonstrate that our method achieves promising quantitative performance and generates natural, human-like motions across a diverse range of HOI tasks.
In summary, our contributions include:
@@
  • We propose the first unified physics-based HOI synthesis framework leveraging the powerful world knowledge of VLMs, supporting long-term human-object interactions with a diverse range of objects, including static, dynamic and articulated ones.
  • We introduce VLM-Guided Relative Movement Dynamics (RMD), a fine-grained spatial-temporal bipartite diagram to automatically construct goal states and reward functions for reinforcement learning. This approach bypasses manual reward engineering while supporting diverse HOI, including both static and dynamic interactions.
@@
  • We introduce Interplay, a novel dataset comprising thousands of interaction plans that cover long-horizon static and dynamic interaction tasks. Extensive simulation results demonstrate the effectiveness of our approach on both single-task and long-horizon multi-task scenarios.
@@ 2 Related Work
Table 1: Comparative analysis of key features between Ours and other methods.
Methods Articulated Object Automated Reward Design Dynamic Object Long-horizon Transition HOI Guidance Granularity Multitask High-level Planning Unified Policy
InterPhys [15] Coarse
InterScene [16] Rule-based FSM Coarse
TokenHSI [17] Rule-based FSM Coarse
UniHSI [13] Coarse
Ours Fine
Human Motion Synthesis with Static Scenes.
@@
In the field of character animation and motion synthesis, most research has focused on human motion interacting with static 3D scenes [18, 19, 20, 21, 22, 23, 24, 25, 26]. Typically, researchers break down complex instructions into several key static poses within the scene [27, 28], and then use motion inpainting and optimization techniques to generate transitions between these key poses [29]. However, this approach often leads to mediocre and inconsistent motions due to the limited expressiveness of the intermediate sequences. Recently, diffusion-based methods like [30, 31, 32, 33, 34, 35] have achieved better results in human-scene synthesis. However, these data-driven approaches are limited to generating short-term motions due to constraints in the dataset, and the physical plausibility of the generated motions is not guaranteed. Apart from data-driven kinematic approaches, some studies have explored the problem within a reinforcement learning framework. For instance,  [36, 37] achieved long-term goal-oriented behaviors by leveraging motion primitives with specific reward designs. The UniHSI [13] framework decomposes instructions into a sequence of point-reaching control tasks under a unified formulation. However, it addresses only the spatial constraints of static scenes while overlooking the temporal dynamics of objects, rendering it inadequate for managing complex dynamic interactions.
@@
Kinematic-based Human Motion Synthesis with Dynamic Objects.
@@
Research in human motion synthesis has increasingly focused on modeling interactions with dynamic objects [38, 39, 5, 40]. Diffusion-based frameworks [41, 42, 43, 44], guide motion generation with object trajectories but often lack realism due to predefined object paths that fail to account for physical plausibility. In contrast, InterDreamer [45] first generates human motion and subsequently uses a pre-trained world model to produce object trajectories, though this approach is limited by the simplicity of the world model, leading to inaccuracies in trajectory prediction. Some studies attempt to address these limitations by jointly modeling human-object interactions with supplementary guidance techniques, including relation intervention [46], contact prediction [47], and affordance estimation [44]. Truman [48] and Lingo [49] trained on a high-quality human-scene interaction (HSI) dataset, achieves dynamic stability through an auto-regressive diffusion model guided by action labels or text. Nevertheless, kinematic models continue to face challenges such as penetration, sliding issues, and difficulties in generating long-term motion, requiring extensive annotations.
@@
Physics-based Human Motion Synthesis with Dynamic Objects.
@@
To generate physically plausible motions, reinforcement learning methods have proven effective in training human-object interaction skills based on motion capture data [4, 50, 51, 52, 3, 53]. AMP [54] introduced an adversarial motion prior framework for realistic motion synthesis. InterPhys [15] further extended this framework by incorporating an HOI motion prior, achieving success in tasks such as sitting, lying, and carrying. Further advancements have led to successful applications in sports activities, including basketball [1, 2], tennis [55], soccer [56], skating [57], and football [58, 59]. TokenHSI [17] integrates multiple skills into a single policy via a task tokenizer. These approaches often depend on heuristic designs and struggle to generalize to longer, multi-round scenarios. Our method seeks to address these limitations by introducing a Relative Movement Dynamics representation, enabling the model to capture both spatial and temporal dynamics effectively. This results in physically plausible and temporally coherent interactions, eliminating the need for manual annotations and enhancing realism in dynamic, long-term HOI tasks. The detailed comparisons of key features are listed in Tab. 1.
@@ 3 Method
Our method comprises two tightly coupled components. First, we describe how a VLM is leveraged to translate high-level instruction and environmental context into a sequence of structured interaction plans. This process is grounded in our proposed concept of RMD, which models fine-grained spatio-temporal relationships between human and object parts and serves as the foundation for interaction planning (Sec. 3.1). Then, we present a VLM-guided motion policy learning framework, where each RMD plan is automatically transformed into goal state and a corresponding reward function, enabling the agent to learn executable motion policies without manual specification (Sec. 3.2).
Figure 2: An overview of our architecture. Receiving instruction and environment context as input, the VLM-Guided RMD Planner generates a multi-step interaction plan in the form of RMD. Based on this plan, our framework automatically designs both goal states and reward functions, enabling the VLM-Guided Motion Policy to execute the interaction step by step.
@@ 3.1 VLM-Guided RMD Planner
We employ GPT-4V [60] as our VLM planner to bridge high-level task instructions and low-level interaction execution. As illustrated in Fig. 2, the planner takes three inputs: a high-level textual instruction I, a top-view contextual image of the environment C, and a set of modular prompts designed to guide the decomposition of the task into detailed interaction plans.
Effectively leveraging the motion imagination capabilities of VLMs for HOI planning necessitates a principled formulation of the interaction process. Simplistic abstractions, such as the chain-of-contacts (CoC) proposed by UniHSI [13], is inadequate: by imposing only transient point-contact constraints that are discarded immediately upon satisfaction, CoC fails to capture the evolving spatio-temporal relationships critical for modeling dynamic and coordinated interactions. In contrast, we propose a novel concept, Relative Movement Dynamics (RMD), which encodes fine-grained spatial and temporal relationships between human body parts and object components throughout the course of an interaction. This structured formulation enables the VLM to reason more effectively about the physical organization of human–object interactions and to produce structured, executable plans aligned with underlying motion dynamics. Our key insight is that human-object interactions can be abstracted as the relative motion between two sets of rigid bodies: human body parts and object parts, evolving over time. We formalize this abstraction as a bipartite graph $\mathcal{B}$ that encodes part-level motion dynamics. Let
$P_{H}=\{p_{h_{1}},p_{h_{2}},\dots,p_{h_{m}}\},\quad P_{O}=\{p_{o_{1}},p_{o_{2}},\dots,p_{o_{n}}\}$ (1)
denote the sets of human and object parts, respectively. Then the graph $\mathcal{B}$ is defined as,
$\mathcal{B}=(V,E,w),\qquad V=P_{H}\cup P_{O},\quad E\subseteq P_{H}\times P_{O},\quad w:E\to\{0,1,2,3\}.$ (2)
Each edge $e_{ij}=(p_{h_{i}},p_{o_{j}})\in E$ connects a human part $p_{h_{i}}\in P_{H}$ to an object part $p_{o_{j}}\in P_{O}$, with an associated weight $w_{ij}\in\{0,1,2,3\}$ that characterizes their relative-motion pattern: $w_{ij}=0$ indicates stationary contact, $w_{ij}=1$ denotes approaching motion, $w_{ij}=2$ represents separating motion, and $w_{ij}=3$ implies no consistent relative trend.
To facilitate the planner’s ability to infer such dynamics, we adopt a modular prompt design strategy. Each prompt segment is crafted to trigger a specific reasoning ability, such as environmental parsing, part-level object understanding, motion dynamic inference, or symbolic representation generation. Through this mechanism, the VLM is guided to produce both the RMD graph $\mathcal{B}$ and two spatial anchors: the target location for the human root $\mathcal{T}_{H}$ and the target location for the object root $\mathcal{T}_{O}$, which together define the global interaction objective.
Finally, the planner outputs an interaction plan $\mathcal{D}$ as a sequence of $N$ structured steps,
$\mathcal{D}=\{\mathcal{G}_{1},\mathcal{G}_{2},\ldots,\mathcal{G}_{N}\},$ (3)
where each step $\mathcal{G}_{i}$ is a triplet that specifies the spatial goals and dynamic pattern,
$\mathcal{G}_{i}=\{\mathcal{T}_{H},\mathcal{T}_{O},\mathcal{B}\}.$ (4)
3.2 Automatic Policy Learning via VLM-Guided RMD
Upon receiving the plan generated by the VLM-guided RMD Planner, the policy produces joint-level torques to actuate the agent and execute the plan. Our physics-based HOI policy learning is formulated within the framework of goal-conditioned reinforcement learning. At each timestep $t$, the agent samples an action from its policy $\pi(a_{t}\mid\mathbf{s}_{t},\mathbf{g}_{t})$, conditioned on the current state $\mathbf{s}_{t}$ and the goal state $\mathbf{g}_{t}$. After executing the action, the environment transitions to the next state $\mathbf{s}_{t+1}$, and the agent receives two types of rewards: a task reward $r^{G}(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1})$, which incentivizes behaviors aligned with achieving the specified goal $\mathbf{g}_{t}$, and a style reward $r^{S}(\mathbf{s}_{t},\mathbf{s}_{t+1})$, which encourages the agent to produce natural behaviors as proposed in AMP [54]. Unlike AMP and most prior methods that rely on manually specifying task-specific goal states and handcrafted reward terms, our framework supports the automatic construction of goal states and reward functions. In the following, we detail how our framework automatically generates both goal states and reward functions from VLM output, enabling the agent to learn behaviors aligned with the intended objectives.
Automatic Goal State Construction via VLM-Guided RMD.
To provide the policy with fine-grained spatio-temporal guidance during human-object interaction, we encode the $i$-th planned interaction step $\mathcal{G}_{i}=\{\mathcal{T}_{H},\mathcal{B},\mathcal{T}_{O}\}$, generated by the VLM, with both kinematic and dynamic information.
For each edge $e_{ij}\in E$, we extract from the simulator the absolute position–velocity pair of the human joint, denoted by $\mathbf{p}_{h_{i}}^{p},\mathbf{p}_{h_{i}}^{v}$, and that of the nearest surface point on the object part, $\mathbf{p}_{o_{j}}^{p},\mathbf{p}_{o_{j}}^{v}$. We then compute their relative quantities in the agent-centric coordinate frame:
$\tilde{\mathbf{p}}_{ij}=\mathbf{p}_{o_{j}}^{p}-\mathbf{p}_{h_{i}}^{p},\quad\tilde{\mathbf{v}}_{ij}=\mathbf{p}_{o_{j}}^{v}-\mathbf{p}_{h_{i}}^{v}.$ (5)
The edge weight $w_{ij}\in\{0,1,2,3\}$ is encoded as a one-hot vector $\mathbf{w}_{ij}^{\prime}\in\{0,1\}^{4}$, and concatenated with the relative position and velocity to form the feature tuple $(\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},\mathbf{w}_{ij}^{\prime})$ at timestep $t$. Stacking these features across all edges yields the RMD state: